The magnetic field in an electromagnetic wave has a peak value given by 177 μT.

What is the peak intensity of this wave?
W/m2

Tries 0/2

What is the average intensity of this wave?
W/m2

Tries 0/2
Answer #1

We know, the average intensity of the wave= (c×B²/2)

here, c = speed of light= 3× 10^8 m/s

B= peak value of magnetic field = 177×10^(-6) T

  =permeability of free space=4π × 10^(-7) weber/amp.m

Now, the average intensity is :

I(avg)=[3×10^8×{177×10^(-6)}²/2×4π×10^(-7)]

=>I(avg) = 3.74×10^6 Watt/m²

Again, peak intensity=I = 2× I(avg)

=> I = 2×3.74×10^6 = 7.48×10^6 Watt/m².

hope your questions have been answered properly.

if it helps please give a like.

happy HomeworkLibing.